geerky42 (geerky42):

Is \(\exists x\in U,~\forall y\in U~~ f(x,y)\) and \(\forall y\in U,~\exists x\in U~~ f(x,y)\) logically equivalent? Does order matter?

OpenStudy (anonymous):

Suppose \(f(x,y)\) is the statement that \(x<y\). To say "\(\exists~x\in U\) such that \(\forall~y\in U\), you have \(x<y\)" means there is a number \(x\) that is smaller than any number \(y\) in the set \(U\). To say "\(\forall~y\in U,~\exists~x\in U\), you have \(x<y\)" means that for any number \(y\) there is something smaller than \(y\). As a counter-example, consider the set of real numbers \(\mathbb{R}\). Let \(x\in\mathbb{R}\). No matter what \(x\) is, \(x\) is not smaller than any other number \(y\). For instance, \(x=0\) is smaller than every \(y>0\), but this is not true if \(y<0\) and thus not true *for any* \(y\in\mathbb{R}\). The second statement is true in this case, if \(U=\mathbb{R}\) then there is always a real number \(x\) smaller than any real number \(y\). To summarize, yes. Order matters for different logical quantifiers.

geerky42 (geerky42):

Very good explanation. Thanks!